OpenStudy (calculusxy):

Evaluate this expression 9(18 x 2) / 4(15 / 5)

OpenStudy (calculusxy):

@Mertsj

OpenStudy (mertsj):

What is 18 x 2?

OpenStudy (calculusxy):

18 times 2

OpenStudy (mertsj):

18 x 2 = 36

OpenStudy (calculusxy):

Can you complete this expression?

OpenStudy (calculusxy):

@Mertsj

OpenStudy (calculusxy):

@Preetha

OpenStudy (jack1):

use bodmas to do equation in correct order then its a simple fraction

OpenStudy (tkhunny):

9(18 x 2) / 4(15 / 5) You have written \(9\cdot \dfrac{18\cdot 2}{4}\cdot \dfrac{15}{5}\). Is that what you intended?